OpenStudy (anonymous):

consider,a gambling game,where you have to throw two dice simultaneously.you would gain Rs2 for any even number that shows up,and lose Re1 for any odd no.that shows up.what is ur expected gain or loss form the game?

OpenStudy (anonymous):

(1/2) Rs2 - (1/2) Re1 or (Rs2 - Re1)/2

OpenStudy (anonymous):

hw do u approach probability questions?

OpenStudy (anonymous):

any insights by meverett

OpenStudy (anonymous):

well the last two questions that you have asked, have been completely different. For your first question, I would use the underline method. If we roll the dice once, we need one fraction _____ If we roll the dice twice, then we need two fractions multiplied together _____ ______ If we roll the dice thrice, then we need three fractions multiplied together _____ ______ _______

OpenStudy (anonymous):

For your second question it is an expected value, so we see how many outcomes...... two ....... win or lose, so two blank lines ______ _______ win lose but we need a probability for each as well as the amount won or loss winning = 1/2 chance since 1/2 the numbers are even or winning = (1/2) Rs2 and losing = - (1/2) Re1 notice the negative sign for losing

OpenStudy (anonymous):

okay......

OpenStudy (anonymous):

each of these questions that you are studying, cover a wide range of probability, probably to become the most successful at probability we will need practice at several of the same type of questions for awhile. Is this a review section that you are working on? A challenge section?

OpenStudy (anonymous):

hey i m working for a entrance exam...can u tell me broadly what sort of ques should i solve